Please help asap!!!Assume you have a hypothesis test as follows. Ho: P1 – P2 = 0 | 1 HA : p1 - P2 # 0 You also know based on two surveys that: S
Answer #1

Solution:

\bar{P} = ( x1 + x2 ) / ( n1 + n2 )

= ( 40 + 12) / ( 90+82)

= 0.30

1 - \bar{P} = 0.70

Test statistics

z = (\hat p1 - \hat p 2 ) / \sqrt{} \bar{P}*(1-\bar{P}) ( 1/n1 + 1/n2 )

= ( 0.45 - 0.15) / \sqrt{} (0.30 * 0.70 ) (1/90 + 1/82 )

= 0.30